# Soft gravitons as Goldstone modes of spontaneously broken asymptotic symmetries in de Sitter

### v1

@@
Soft gravitons as Goldstone modes of spontaneously broken asymptotic symmetries in de Sitter
Martin S. Sloth111sloth@sdu.dk
Universe-Origins, University of Southern Denmark,
Campusvej 55, 5230 Odense M, Denmark
@@
I demonstrate that soft graviton modes in de Sitter are the Goldstone modes of the spontaneously broken asymptotic symmetry group of de Sitter. I then show that any local measurement, including the effects of the environment, will collapse the symmetric state onto the broken state in the large volume limit. In any discussion involving observers, de Sitter is, therefore, best described globally by the broken phase, while local observers, in the small volume limit, can not discriminate between different degenerate global vacuum states, and are therefore best described by the symmetric state. This illuminates the physical nature of soft graviton modes in de Sitter.
1 Introduction
@@
There has been a long history of discussions of the physical reality of infrared super-horizon (soft) graviton modes in de Sitter among practitioners of cosmology. The opinion ranges from arguments that they are pure gauge [1, 2, 3, 4], with no physical meaning, to arguments that they are responsible for the ultimate demise of de Sitter by the decay of the cosmological constant [5, 7, 6, 8, 10, 9, 11, 12, 13, 14, 15] . In between, we find the more moderate point of view that they are a pure gauge locally but do have global effects once the gauge invariant observers are properly defined and included as part of the quantum system. The effects can be observed in the Cosmic Microwave Background (CMB) or Large Scale Structure of the universe today when modes enter the horizon after the end of the quasi-de Sitter inflationary phase in the early universe or as deformations of the global properties of de Sitter [16, 17, 18, 19, 20, 21]. Here I am going to further clarify the physical understanding of soft graviton modes by showing that they can be viewed as the Goldstone modes of spontaneously broken asymptotic symmetries in de Sitter.
@@
It was argued some years ago in the context of black holes that the Goldstone modes of the spontaneously broken asymptotic symmetries of the black holes, the BMS supertranslations, are the soft graviton modes satisfying Weinberg’s soft graviton theorems [22, 23]. There is a very analogous story in de Sitter space. The equivalent of the Weinberg soft theorem satisfied by the Weinberg adiabatic soft graviton modes is the Maldacena consistency relations, and the asymptotic symmetries of de Sitter are the diffeomorphisms in three-dimensional Euclidian space [24]. Here I will show that the soft graviton modes are, indeed, also the Goldstone modes of the spontaneously broken asymptotic symmetry group of de Sitter. In cosmology, we also think of soft graviton modes in de Sitter, as frozen-out super-horizon gravitational fluctuations, described by the traceless and transverse spatial component of the graviton.
@@
In order to show that the super-horizon soft graviton is the Goldstone of the spontaneously broken asymptotic symmetry group of de Sitter, we will discuss the symmetry-breaking pattern and then show that the counting of Goldstone modes adds up correctly. At the linearized level, the anisotropic scaling symmetry generated by symmetric traceless matrices is broken, leaving five broken generators. However, in the case of broken space-time symmetries, it is well-known that Goldstone’s theorem for internal symmetries does not trivially apply. As in the case with spontaneously broken non-Lorentz invariant symmetries [25], also in the case of space-time symmetries, the number of massless Goldstone modes can be different than the number of broken symmetry generators. In the case of unbroken translations, local translation in the unbroken directions can be used to relate the Goldstone modes to each other [26]. With three unbroken translations only two physical Goldstone modes remain out of the five broken generators. As we will show below, these are the two massless Goldstone modes of the soft graviton, which remain as physical modes. This argument generalizes to any dimension.
@@
In the next section, section 2, I discuss the symmetry-breaking pattern and how to count the Goldstone modes correctly. Then, in section 3, I discuss the role of local observers versus the role of global observers. I briefly discuss the generalization of these arguments to a period of slow-roll inflation where the time-translation invariance of de Sitter is softly broken in section 4. Then, in section 5, I discuss some examples from the literature, and how they relate to the interpretation of soft gravitons as the Goldstone modes of spontaneously broken asymptotic symmetries. Finally, I conclude in section 6.
2 Spontaneously broken asymptotic symmetries and soft modes as Goldstones
@@
The solutions of Einstein’s equations with a cosmological constant break the full diffeomorphism symmetry down to the isometry group of de Sitter, $SO(D,1)$ in $D=1+n$ spacetime dimensions. However, an observer living on the space-like infinity of de Sitter would experience a higher symmetry and conclude that the $SO(D,1)$ de Sitter invariance is enhanced to the asymptotic symmetry group of de Sitter, which is the diffeomorphisms of $\mathbb{R}^{n}$ [24].
@@
In $n$ Euclidian dimensions the general covariance implies that the index, $j$, up or down transforms as covariant or contravariant vectors under the linear matrix multiplication of matrices given by $\partial x^{i}/\partial\tilde{x}^{j}$ $\partial\tilde{x}^{j}/\partial x^{i}$ at the level of linearized infinitesimal coordinate transformation. The linearized coordinate transformations are elements of the symmetry group $GL(n,\mathbb{R})$ with dimension $n^{2}$. $GL(n,\mathbb{R})$ can be written as $SL(n,\mathbb{R})\times\mathbb{R}$, where $SL(n,\mathbb{R})$ has $SO(n)$ as a subgroup with dimension $n(n-1)/2$ generated by the antisymmetric tensors $J_{ij}$, while the other $n(n+1)/2-1$ symmetric traceless generators $S_{ij}$ generate shear deformations. The dilatation, $D$, is part of $GL(n,\mathbb{R})$, but not $SL(n,\mathbb{R})$, which has dimension $n^{2}-1$. Including translation invariance enhances $GL(n,\mathbb{R})$ to the affine group $Aff(n,R))=GL(n,\mathbb{R})\times P_{n}$, where $P_{n}$ is the group of translations. The affine group, therefore, has the Poincare group as a subgroup. The full diffeomorphism group is the closure of the affine group and the special conformal group. Any generator of the full diffeomorphism group can be written in terms of (infinite) combinations of the generators of $SL(n,\mathbb{R})$, as well as the generators of translations $P_{n}$, dilatations $D$ and special conformal transformations $K_{n}$ [27]. Important for the discussion below, at the linearized level, the generators of the asymptotic symmetry group are therefore $S_{ij},J_{ij},P_{i},K_{i},D$.
@@
When we restrict ourselves to pure de Sitter (with unbroken time-translation), the relevant gravitational degrees of freedom are contained in the transverse and traceless graviton $\gamma_{ij}$, which be extracted from the metric in the planar slicing
$ds^{2}=-dt^{2}+a^{2}(t)\left[e^{\gamma}\right]_{ij}dx^{i}dx^{j}\,.$ (2.1)
@@
We have ignored the lapse and shift, which decay, and therefore are unimportant on the asymptotic boundary. The asymptotic symmetry transformations act on this metric as [28]
$\delta\left(\left[e^{\gamma}\right]_{ij}\right)={\cal L}_{\xi}\left(\left[e^{\gamma}\right]_{ij}\right)~{}.$ (2.2)
@@
$\delta\gamma_{ij}=\partial_{i}\xi_{j}+\partial_{j}\xi_{i}-\frac{2}{3}\partial\cdot\xi\delta_{ij}+\mathcal{O}(\gamma).$ (2.3)
@@
$\partial_{i}\xi_{j}+\partial_{j}\xi_{i}-\frac{2}{3}\partial\cdot\xi\delta_{ij}=0$ (2.4)
can not generate tensor adiabatic modes. Eq. (2.4) is solved by
$\xi_{i}=c_{i}+\omega_{ij}x_{j}+cx_{i}+\Big{(}x^{2}a_{i}-2x_{i}(\vec{x}\cdot\vec{a})\Big{)}$ (2.5)
where $c_{i}$, $a_{i}$, $c$ and $\omega_{ij}=-\omega_{ji}$ is constant parameters. One recognizes (2.5) as the conformal Killing vector of $\mathbb{R}^{3}$, where $c_{i}$ is associated to the translations, $c$ to dilations, $a_{i}$ to conformal boosts and $\omega_{ij}$ to rotations.
@@
$S_{ij}=x_{i}\partial_{j}+x_{j}\partial_{i}-\frac{2}{3}\delta_{ij}x^{k}\partial_{k}$ (2.6)
@@
$\xi_{i}^{L}=-\frac{1}{2}\gamma_{ij}^{L}x^{j}$ (2.9)
implying $\delta\gamma_{ij}=-\gamma_{ij}^{L}$. This amounts to switching on a non-zero vacuum expectation value for the graviton $\gamma^{L}_{ij}=\left<\gamma_{ij}(x)\right>$.
Therefore, in four space-time dimensions with $n=3$, we have that at the linearized level, the soft graviton mode is breaking $SL(3)$ down to $SO(3)$ and so one would expect a ${\rm dim}\,SL(3)/SO(3)=5$ Goldstone modes, corresponding to the five broken generators, $S_{ij}$, which is inconsistent with the number of physical graviton modes. The reason is that one may use the diffeomorphism group to cancel some of the 5 degrees of freedom of the adiabatic mode. It was shown in [26] that in the directions of unbroken spatial translations, one may use a local translation to remove the redundant degrees of freedom. This is only true when the commutator of the generators of the broken symmetries does not commute with the unbroken translation generators. That is why for internal symmetries where the generators of broken internal symmetries commute with the translations, this does not affect the counting of Goldstone modes, while in the case of broken space-time symmetries, which do not commute with the unbroken translation generators, the number of Goldstone modes is less than the number of broken generators. Indeed, an $x$-depended translation is generated by
$\xi_{i}=c_{i}(x)$ (2.10)
@@
$\delta\gamma_{ij}=\partial_{i}c_{j}(x)+\partial_{j}c_{i}(x)-\frac{2}{3}\delta_{ij}\partial_{k}c_{k}(x).$ (2.11)
Clearly now, we can cancel part of (2.7) by choosing
$\partial_{i}c_{j}(x)+\partial_{j}c_{i}(x)-\frac{2}{3}\delta_{ij}\partial_{k}c_{k}(x)=-s_{ij}$ (2.12)
@@
Following [19] we can extend the transformation in eq.(2.3) to higher order in $\gamma$. To next order in $\gamma$, we obtain
$\delta\gamma_{ij}=\partial_{i}\xi_{j}+\partial_{j}\xi_{i}-\frac{2}{3}\partial\cdot\xi\delta_{ij}+\xi\cdot\partial\gamma_{ij}+\mathcal{O}(\gamma^{2})\,.$ (2.13)
In general, for a spontaneously broken symmetry generated by the charge, $Q$, there exists an order parameter $[Q,\mathcal{O}]=-i\delta\mathcal{O}$ for some operator $\mathcal{O}$ [29]. Using now $\xi=\xi_{i}^{L}$ given by eq.(2.9) in eq.(2.13) , and substituting the expression into the standard expression for the Nöther charge
$Q(\xi)=\frac{1}{2}\int d^{3}x\Pi_{\gamma}^{ij}\delta\gamma_{ij}~{}+~{}h.c.~{},$ (2.14)
@@
$\langle 0|[Q,\gamma_{ij}]|0\rangle=i\gamma_{ij}^{L}\,,$ (2.15)
by using the standard canonical commutation relations. This is as expected when $Q$ is the charge generating a spontaneously broken symmetry transformation and $\gamma_{ij}^{L}$ is the order parameter, and we conclude that $Q$ generates the shift $\delta\gamma_{ij}=-\gamma_{ij}^{L}$ in $\gamma_{ij}$ [19].
We saw above that rotational invariance is unbroken at zeroth order in $\gamma$. However, since the order parameter is not invariant under rotations, it breaks rotational symmetry. Rotational symmetry is, therefore, explicitly broken at linear order in $\gamma$. In fact, this can be seen from taking $\xi_{i}=-\omega_{ij}x^{j}$, with $\omega_{ij}=-\omega_{ji}$ corresponding to a rotation. In this case, we have, when the rotation acts on $\gamma_{ij}$ at higher order, $\delta\gamma_{ij}=\omega_{jl}\gamma_{li}+\omega_{il}\gamma_{lj}$, and so $\partial_{i}\delta\gamma_{ij}=\partial_{i}\omega_{il}\gamma_{lj}\neq 0$, such that $\delta\gamma_{ij}$ generated by a rotation acting on $\gamma_{ij}$ is not transverse, meaning that the rotations leave the transverse and traceless gauge at linear order in $\gamma$. Thus at second order, rotational symmetry is explicitly broken by the presence of the order parameter. By using that the angular momentum operator takes the form $J^{i}\propto\int d^{x}\epsilon^{ijk}\gamma_{mk}\dot{\gamma}_{ml}$, one also finds $[J^{i},\gamma_{ij}]\neq 0$, as expected for the generator of a broken symmetry. So, while the anisotropic scaling symmetries, the shear transformations, are spontaneously broken at zeroth order in $\gamma$, their breaking induces an explicit breaking of rotational symmetry at higher order.
We have found that the order parameter is a symmetric and traceless tensor, which breaks rotational invariance, like in the case of a Nematic liquid crystal, although, the spontaneous symmetry breaking pattern is different. In Nematic liquid crystals, the rotational symmetry is spontaneously broken, while anisotropic scaling is explicitly broken. Here it is the opposite.
The Goldstone modes are the long-range spatial fluctuations of the order parameter $\gamma_{ij}^{L}$ Thus, the soft modes are Goldstones of the spontaneously broken asymptotic symmetries. This also means that the Lagrangian for the Goldstone modes are of the general form [30]
$\mathcal{L}(\pi)\propto g_{ab}\partial_{\mu}\pi^{a}\partial^{\mu}\pi^{b}+\mathcal{O}(\partial^{4}),$ (2.16)
@@
$\mathcal{L}(\gamma)\propto-\dot{\gamma}_{ij}(x)\dot{\gamma}_{ij}(x)+a^{-2}\partial_{l}\gamma_{ij}(x)\partial_{l}\gamma_{ij}(x)~{},$ (2.17)
which happens to agree with the quadratic action of (soft) gravitons in de Sitter.
3 Symmetry breaking: local versus global observers

### v2

@@
Abstract
@@
I demonstrate that soft graviton modes in de Sitter spacetimes are the Goldstone modes of the spontaneously broken asymptotic symmetry group of de Sitter space. I then show that any local measurement, including the effects of the environment, will collapse the symmetric state onto the broken state in the large volume limit. In any discussion involving observers, de Sitter spacetimes are, therefore, best described globally by the broken phase, while local observers, in the small volume limit, can not discriminate between different degenerate global vacuum states and are therefore best described by the symmetric state. As a consequence, a small Hubbled-sized local region initially in the symmetric state will, after a time-scale corresponding to the Page time of de Sitter space, have expanded to a large region in the broken state. This illuminates the physical nature of soft graviton modes in de Sitter spacetimes.
I Introduction
@@
There has been a long history of discussions of the physical reality of infrared super-horizon (soft) graviton modes in de Sitter spacetimes among practitioners of cosmology. The opinion ranges from arguments that they are pure gauge Allen:1986ta ; Allen:1986tt ; Higuchi:2000ye ; Higuchi:2011vw , with no physical meaning, to arguments that they are responsible for the ultimate demise of de Sitter space by the decay of the cosmological constant Polyakov:1982ug ; Polyakov:2009nq ; Polyakov:2007mm ; Tsamis:1992sx ; Tsamis:1994ca ; Tsamis:1993ub ; Tsamis:2007is ; Myhrvold:1983hu ; Dvali:2013eja ; Dvali:2014gua ; Dvali:2018jhn . In between, we find the more moderate point of view that they are a pure gauge locally but do have global effects once the gauge invariant observers are properly defined and included as part of the quantum system. The effects can be observed in the cosmic microwave background or large scale structure of the universe today when modes enter the horizon after the end of the quasi–de Sitter inflationary phase in the early universe or as deformations of the global properties of de Sitter spacetimes Giddings:2011zd ; Giddings:2011ze ; Giddings:2010nc ; Ferreira:2016hee ; Ferreira:2017ogo ; Ferreira:2017erz . Here I am going to further clarify the physical understanding of soft graviton modes by showing that they can be viewed as the Goldstone modes of spontaneously broken asymptotic symmetries in de Sitter space.
@@
It was argued some years ago in the context of black holes that the Goldstone modes of the spontaneously broken asymptotic symmetries of the black holes, the Bondi-Metzner-Sachs (BMS) supertranslations, are the soft graviton modes satisfying Weinberg’s soft graviton theorems He:2014laa ; Hawking:2016msc . There is a very analogous story in de Sitter space. The equivalent of the Weinberg soft theorem satisfied by the Weinberg adiabatic soft graviton modes is the Maldacena consistency relations, and the asymptotic symmetries of de Sitter space are the diffeomorphisms in three-dimensional Euclidean space Anninos:2010zf . Here I will show that the soft graviton modes are, indeed, also the Goldstone modes of the spontaneously broken asymptotic symmetry group of de Sitter space. In cosmology, we also think of soft graviton modes in de Sitter spacetimes as frozen-out super-horizon gravitational fluctuations, described by the traceless and transverse spatial component of the graviton.
@@
In order to show that the superhorizon soft graviton is the Goldstone of the spontaneously broken asymptotic symmetry group of de Sitter space, I will discuss the symmetry-breaking pattern and then show that the counting of Goldstone modes adds up correctly. At the linearized level, the anisotropic scaling symmetry generated by symmetric traceless matrices is broken, leaving five broken generators. However, in the case of broken space-time symmetries, it is well known that Goldstone’s theorem for internal symmetries does not trivially apply. As in the case with spontaneously broken non-Lorentz invariant symmetries Nielsen:1975hm , also in the case of space-time symmetries, the number of massless Goldstone modes can be different than the number of broken symmetry generators. In the case of unbroken translations, local translation in the unbroken directions can be used to relate the Goldstone modes to each other Ivanov:1975zq ; Low:2001bw . With three unbroken translations, only two physical Goldstone modes remain out of the five broken generators. As I will show below, these are the two massless Goldstone modes of the soft graviton, which remain as physical modes. This argument generalizes to any dimension.
@@
In the next section, I discuss the symmetry-breaking pattern and how to count the Goldstone modes correctly. Then I discuss the role of local observers versus the role of global observers. Afterwards, I briefly discuss the generalization of these arguments to a period of slow-roll inflation where the time-translation invariance of de Sitter space is softly broken. Then, I discuss some examples from the literature and how they relate to the interpretation of soft gravitons as the Goldstone modes of spontaneously broken asymptotic symmetries. Finally, I conclude.
II Spontaneously broken asymptotic symmetries and soft modes as Goldstones
@@
The solutions of Einstein’s equations with a cosmological constant break the full diffeomorphism symmetry down to the isometry group of de Sitter space, $SO(D,1)$ in $D=1+n$ spacetime dimensions. However, an observer living on the space-like infinity of de Sitter space would experience a higher symmetry and conclude that the $SO(D,1)$ de Sitter invariance is enhanced to the asymptotic symmetry group of de Sitter space, which is the diffeomorphisms of ${\mathbb{R}}^{n}$ Anninos:2010zf .
@@
In $n$ Euclidian dimensions the general covariance implies that the index, $j$, up or down transforms as covariant or contravariant vectors under the linear matrix multiplication of matrices given by $\partial x^{i}/\partial\tilde{x}^{j}$ $\partial\tilde{x}^{j}/\partial x^{i}$ at the level of linearized infinitesimal coordinate transformation. The linearized coordinate transformations are elements of the symmetry group $GL(n,\mathbb{R})$ with dimension $n^{2}$. $GL(n,\mathbb{R})$ can be written as $SL(n,\mathbb{R})\times\mathbb{R}$, where $SL(n,\mathbb{R})$ has $SO(n)$ as a subgroup with dimension $n(n-1)/2$ generated by the antisymmetric tensors $J_{ij}$, while the other $n(n+1)/2-1$ symmetric traceless generators $S_{ij}$ generate shear deformations. The dilatation, $D$, is part of $GL(n,\mathbb{R})$, but not $SL(n,\mathbb{R})$, which has dimension $n^{2}-1$. Including translation invariance enhances $GL(n,\mathbb{R})$ to the affine group $Aff(n,R))=GL(n,\mathbb{R})\times P_{n}$, where $P_{n}$ is the group of translations. The affine group, therefore, has the Poincaré group as a subgroup. The full diffeomorphism group is the closure of the affine group and the special conformal group. Any generator of the full diffeomorphism group can be written in terms of (infinite) combinations of the generators of $SL(n,\mathbb{R})$, as well as the generators of translations $P_{n}$, dilatations $D$ and special conformal transformations $K_{n}$ Borisov:1974bn . Important for the discussion below, at the linearized level, the generators of the asymptotic symmetry group are therefore $S_{ij},J_{ij},P_{i},K_{i},D$.
@@
When we restrict ourselves to pure de Sitter space (with unbroken time translation), the relevant gravitational degrees of freedom are contained in the transverse and traceless graviton $\gamma_{ij}$, which be extracted from the metric in the planar slicing
$ds^{2}=-dt^{2}+a^{2}(t)\left[e^{\gamma}\right]_{ij}dx^{i}dx^{j}\,.$ (1)
@@
We have ignored the lapse and shift, which decay, and therefore are unimportant on the asymptotic boundary. The asymptotic symmetry transformations act on this metric as 1203.6351
$\delta\left(\left[e^{\gamma}\right]_{ij}\right)={\cal L}_{\xi}\left(\left[e^{\gamma}\right]_{ij}\right)~{}.$ (2)
@@
$\delta\gamma_{ij}=\partial_{i}\xi_{j}+\partial_{j}\xi_{i}-\frac{2}{3}\partial\cdot\xi\delta_{ij}+\mathcal{O}(\gamma).$ (3)
@@
$\partial_{i}\xi_{j}+\partial_{j}\xi_{i}-\frac{2}{3}\partial\cdot\xi\delta_{ij}=0$ (4)
can not generate tensor adiabatic modes. Equation (4) is solved by
$\xi_{i}=c_{i}+\omega_{ij}x_{j}+cx_{i}+\Big{(}x^{2}a_{i}-2x_{i}(\vec{x}\cdot\vec{a})\Big{)}$ (5)
where $c_{i}$, $a_{i}$, $c$, and $\omega_{ij}=-\omega_{ji}$ are constant parameters. One recognizes (5) as the conformal Killing vector of $\mathbb{R}^{3}$, where $c_{i}$ is associated to the translations, $c$ to dilations, $a_{i}$ to conformal boosts and $\omega_{ij}$ to rotations.
@@
$S_{ij}=x_{i}\partial_{j}+x_{j}\partial_{i}-\frac{2}{3}\delta_{ij}x^{k}\partial_{k}$ (6)
@@
$\xi_{i}^{L}=-\frac{1}{2}\gamma_{ij}^{L}x^{j}$ (9)
implying $\delta\gamma_{ij}=-\gamma_{ij}^{L}$. This amounts to switching on a nonzero vacuum expectation value for the graviton $\gamma^{L}_{ij}=\left<\gamma_{ij}(x)\right>$.
Therefore, in four space-time dimensions with $n=3$, we have that at the linearized level, the soft graviton mode is breaking $SL(3)$ down to $SO(3)$ and so one would expect a ${\rm dim}\,SL(3)/SO(3)=5$ Goldstone modes, corresponding to the five broken generators, $S_{ij}$, which is inconsistent with the number of physical graviton modes. The reason is that one may use the diffeomorphism group to cancel some of the five degrees of freedom of the adiabatic mode. It was shown in Ivanov:1975zq ; Low:2001bw that in the directions of unbroken spatial translations, one may use a local translation to remove the redundant degrees of freedom. This is only true when the commutator of the generators of the broken symmetries does not commute with the unbroken translation generators. That is why for internal symmetries where the generators of broken internal symmetries commute with the translations, this does not affect the counting of Goldstone modes, while in the case of broken space-time symmetries, which do not commute with the unbroken translation generators, the number of Goldstone modes is less than the number of broken generators. Indeed, an $x$-dependent translation is generated by
$\xi_{i}=c_{i}(x)$ (10)
@@
$\delta\gamma_{ij}=\partial_{i}c_{j}(x)+\partial_{j}c_{i}(x)-\frac{2}{3}\delta_{ij}\partial_{k}c_{k}(x).$ (11)
Clearly now, we can cancel part of (7) by choosing
$\partial_{i}c_{j}(x)+\partial_{j}c_{i}(x)-\frac{2}{3}\delta_{ij}\partial_{k}c_{k}(x)=-s_{ij}$ (12)
@@
Following Ferreira:2016hee , we can extend the transformation in Eq.(3) to higher order in $\gamma$. To next order in $\gamma$, we obtain
$\delta\gamma_{ij}=\partial_{i}\xi_{j}+\partial_{j}\xi_{i}-\frac{2}{3}\partial\cdot\xi\delta_{ij}+\xi\cdot\partial\gamma_{ij}+\mathcal{O}(\gamma^{2})\,.$ (13)
In general, for a spontaneously broken symmetry generated by the charge, $Q$, there exists an order parameter $[Q,\mathcal{O}]=-i\delta\mathcal{O}$ for some operator $\mathcal{O}$ Brauner:2010wm . Using now $\xi=\xi_{i}^{L}$ given by Eq.(9) in Eq.(13), and substituting the expression into the standard expression for the Nöther charge
$Q(\xi)=\frac{1}{2}\int d^{3}x\Pi_{\gamma}^{ij}\delta\gamma_{ij}~{}+~{}h.c.~{},$ (14)
@@
$\langle 0|[Q,\gamma_{ij}]|0\rangle=i\gamma_{ij}^{L}\,,$ (15)
by using the standard canonical commutation relations. This is as expected when $Q$ is the charge generating a spontaneously broken symmetry transformation and $\gamma_{ij}^{L}$ is the order parameter, and we conclude that $Q$ generates the shift $\delta\gamma_{ij}=-\gamma_{ij}^{L}$ in $\gamma_{ij}$ Ferreira:2016hee .
We saw above that rotational invariance is unbroken at zeroth order in $\gamma$. However, since the order parameter is not invariant under rotations, it breaks rotational symmetry. Rotational symmetry is, therefore, broken at linear order in $\gamma$. In fact, this can be seen from taking $\xi_{i}=-\omega_{ij}x^{j}$, with $\omega_{ij}=-\omega_{ji}$ corresponding to a rotation. In this case, we have, when the rotation acts on $\gamma_{ij}$ at higher order, $\delta\gamma_{ij}=\omega_{jl}\gamma_{li}+\omega_{il}\gamma_{lj}$, and so $\partial_{i}\delta\gamma_{ij}=\partial_{i}\omega_{il}\gamma_{lj}\neq 0$, such that $\delta\gamma_{ij}$ generated by a rotation acting on $\gamma_{ij}(x)$ is not transverse, meaning that the rotations leave the transverse and traceless gauge at linear order in $\gamma$. Thus, at second order, rotational symmetry is nonlinearly realized on the Goldstone mode. By using that the angular momentum operator takes the form $J^{i}\propto\int d^{3}x\epsilon^{ijk}\gamma_{mk}\dot{\gamma}_{ml}$, one also finds $[J^{i},\gamma_{ij}]\neq 0$, as expected for the generator of a broken symmetry. So, while the anisotropic scaling symmetries, the shear transformations, are spontaneously broken at zeroth order in $\gamma$, their breaking induces a breaking of rotational symmetry at higher order.
We have found that the order parameter is a symmetric and traceless tensor, which breaks rotational invariance, like in the case of a Nematic liquid crystal, although the spontaneous symmetry-breaking pattern is different. In Nematic liquid crystals, the rotational symmetry is spontaneously broken, while anisotropic scaling is explicitly broken.
The Goldstone modes are the long-range spatial fluctuations of the order parameter $\gamma_{ij}^{L}$ Thus, the soft modes are Goldstones of the spontaneously broken asymptotic symmetries. This also means that the Lagrangian for the Goldstone modes is of the general form Naegels:2021ivf
$\mathcal{L}(\pi)\propto g_{ab}\partial_{\mu}\pi^{a}\partial^{\mu}\pi^{b}+\mathcal{O}(\partial^{4}),$ (16)
@@
$\mathcal{L}(\gamma)\propto-\dot{\gamma}_{ij}(x)\dot{\gamma}_{ij}(x)+a^{-2}\partial_{l}\gamma_{ij}(x)\partial_{l}\gamma_{ij}(x)~{},$ (17)
which happens to agree with the quadratic action of (soft) gravitons in de Sitter spacetimes.
III Symmetry breaking: local versus global observers